### v1

# The exact convex roof for GHZ-W mixtures for three qubits and beyond

Abstract
I present an exact solution for the convex roof of the square root threetangle for all states within the Bloch sphere. The working hypothesis is that optimal decompositions contain as many states from the zero-polytope as possible which can be called zero-state locking. The footprint of the measure of entanglement consists in a characteristic pattern for the fixed pure states on the surface which form the optimal solution. The solution is subject to transformation properties due to the SL-invariance of the entanglement measure.
I Introduction
Although we know that entanglement is everywhere present where there is interaction, it is difficult to measure it, because the operator one has to measure is non-Hermitean and consequently the eigenvalues are complex values for every pure states. We therefore take its absolute value to quantify this important resource, which is indispensable for quantum computingLadd et al. (2010); Gill et al. (2022); Orús et al. (2019); Ringbauer et al. (2022), quantum communicationGisin and Thew (2007); Luo et al. (2023); Cozzolino et al. (2019); De Santis et al. (2024), and quantum sensing in generalNarducci et al. (2022); Degen et al. (2017); Crawford et al. (2021); Wang et al. (2023); Aslam et al. (2023); Schnabel et al. (2010); Pezze et al. (2018); Polino et al. (2020); quantum technology is working because these non-local quantum correlations are there. Reasonably, as a non-local quantity, entanglement has to be invariant under the change of the local basis and permutations of the local entities. This leads to the concept of LOCC invariance and the monotone propertyVidal (2000) that every good entanglement measure must satisfy. In general one has to prove this monotone property, which is granted for, when extending the local invariance group from $SU(n)$ to $SL(n)$Verstraete et al. (2003) for $n$ being the number of different axis of the local entity one can do rotations about. In general, the quantities obtained through this extension have the elegant property of being merely susceptible for entanglement at a global level, discarding everything which is non-global. Thus, they are measures of genuine multipartite SL-entanglementOsterloh and Siewert (2005); Doković and Osterloh (2009). Examples for taking qubits as local entities are the concurrenceHill and Wootters (1997); Wootters (1998) and the threetangleCoffman et al. (2000).
The monotone property has included already the convex-roof extension to arbitrary mixed states. However, the calculation of it turns out to be a difficult task, which even is thought to be NP-hardGurvits (2003, 2004) (see however Wen et al. (2024)). For very few examples we have an exact solution for evaluating this convex-roof, as for the concurrence for two qubits. These are examples for which the pure state measure scales linear in $\rho$; it generally holds if the corresponding $d$-th root of a tangle of degree $2d$ is connected to a bilinear functional as does the concurrence. This happens for the threetangle but only for certain statesNeveling and Osterloh (2024). This is not so in general, and the zero-polytope consists of four independent solutions. However, for the mixture of GHZ and W states and the square root of the threetangle it is knownLohmayer et al. (2006); Eltschka et al. (2008); Viehmann et al. (2012) that the Bloch-sphere has two tetrahedra in it: one is the zero-polytope and the other is the polytope made of three of the states of the zero-polytope and the GHZ state.
Based on the conjectured property of zero-state lockingOsterloh (2016); Neveling and Osterloh (2024), in a recent workNeveling and Osterloh (2024) it has been established how optimal decompositions will behave for rank-$2$ density matrices (see Fig. 1). Optimal decompositions are limited to at most $4$ statesCarathéodory (1911); Uhlmann (1998), i.e. to three dimensional simplices for this case.
Figure 1: Polytopes of type $3^{+}N$ (left figure) and $3^{+}S$ (right figure), taken from Ref. Neveling and Osterloh (2024). To the zero polytope (blue) entangled polytopes (gray) are added with corresponding pure states $|N_{\pm}\rangle$. It was noted that a pure state will be added for every of the faces of the zero polytope, which are four in this particular case. For symmetry requirements $|N_{\pm}\rangle$, as well as the pure states of the $(2,1)$ decompositions, are situated in $\phi=0$. This condition will be relaxed with the symmetry here, because of the non-intersection property of optimal decompositions. The optimal decompostions, here indicated by green lines, brachiate through this structure of polytopes with the additional pure state at $\phi=0$.
We use the nomenclature $(n_{z},n_{e})$ of this work for identifying the decompositions, where $n_{z}$ is the number of states from the zero polytope, and $n_{e}$ that of entangled pure states in the decomposition. In particular, it has been observed that for each three surface elements of the zero polytope there is precisely one entangled pure state $|N\rangle$ forming a $(3,1)$ polytope. Within polytopes the entanglement varies linearly due to convexification. These pure states $|N_{i}\rangle$ need to be found as a convexification point of $(1,1)$ and $(2,1)$ decompositions. Optimal decompositions do not intersect, and this fills out good part of the Bloch-sphere. Then, for the tips of each two connected surface elements, $|N_{i_{1}}\rangle$ and $|N_{i_{2}}\rangle$, out of the pure entangled states $|N_{i}\rangle$ found before, are connected by a line of pure entangled states. These give further $(2,1)$ decompositions in the Bloch sphere. All remaining decompositions will be of the type $(1,1)$. This excludes the possibility of optimal $(0,n_{e})$ decompositions for $n_{e}>1$; there are strong arguments for this to holdNeveling and Osterloh (2024), but no general proof exists. Optimality of $(n_{z},n_{e})$ have been excluded numerically through case studiesNeveling and Osterloh (2024). The transformation property, also of optimal decomposition statesViehmann et al. (2011, 2012), is an important finding and lifts the results of this work to all states being SL-equivalent to the mixed state in consideration.
This work is laid out as follows. At first we will deduce those pure states $|N_{i}\rangle$ which form optimal $(3,1)$ decompositions, then find the lines interconnecting each pair of the entangled pure states $|N_{i}\rangle$. The remaining space of the Bloch-sphere is filled with $(1,1)$ decompositions.
The pure entangled states $|N_{i}\rangle$ We analyze the Bloch sphere of the rank-two mixed state
$\rho[p]=p|GHZ\rangle\langle GHZ|+(1-p)|W\rangle\langle W|$ (1)
where $|GHZ\rangle=(|111\rangle+|000\rangle)/\sqrt{2}$ and $|W\rangle=(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$. A three-fold rotational symmetry due to cyclic permutations of the qubits exist for both states. Therefore, it is enough to perform the classification in a wedge of the Bloch sphere given by the azimuthal angle $\phi\in[-\pi/3,\pi/3]$ only. The rest of the sphere is obtained by symmetry requirements.
One of the pure entangled state is the maximally entangled GHZ stateLohmayer et al. (2006); Viehmann et al. (2012). This is also confirmed through the convexification procedure of $(2,1)$ and $(1,1)$ decompositions which correctly points to the GHZ state. Remains to find the state in one of the three vertical wedges of the Bloch sphere. One possibility is to choose another orthogonal basis for the same Bloch sphere. The one orthogonal decomposition chosen here is the real one of the three pure states of the zero-polytope together with its orthogonal partner. These are given by
$|Z_{3}\rangle$ $=$ $\sqrt{p_{0}}|GHZ\rangle-\sqrt{1-p_{0}}|W\rangle$ (2)
$|Z_{3;\perp}\rangle$ $=$ $\sqrt{1-p_{0}}|GHZ\rangle+\sqrt{p_{0}}|W\rangle$ (3)
[…]
Figure 2: Convexification procedure of $|Z3\rangle$ mixed with the corresponding orthogonal state. The curves are made convex with a line touching at $p_{c}=0.0964142$, as is highlighted in the inset, were we plot the differenz in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line. The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are shown in blue above both curves.
It shows that at $p_{c}=0.0964142$ the line is a proper convexification, starting from zero and touching the curve for $(2,1)$ decompositions; the $(1,1)$ decompositions are shown in blue. The corresponding pure state is located in $p_{2}(p_{1},\phi_{1},p_{c})=0.00673174$. Since $p_{c}<1/2$, this corresponds to an angle $\theta_{1}=\pi-\arccos{(2p_{c}-1)}=0.164279=9.4125^{\circ}$, which adds up to the angle of $|Z_{3;\perp}\rangle$, which is $\theta_{0}=1.8273=104.697^{\circ}$. The angle of the new state is hence calculated to be $\theta=\theta_{0}+\theta_{1}=1.99158=114.109^{\circ}$. The four tetrahedra are shown in grey in Fig. 7 with green edges. This is confirmed from the convexification procedure with the $(1,1)$ decompositions (with $|W\rangle$ here) for eigenstates being situated at an angle $\theta=140^{\circ}$. Invoking the three-fold symmetry for this case leads to the three optimal $(3,1)$ decompositions with the states $|N_{i}\rangle$, $i=1,2,3$, as a tip. This shows as well that in this case $(2,1)$ decompositions are optimal in the direction of $|GHZ\rangle$ and $(1,1)$ decompositions are optimal in the direction of $|W\rangle$. Fig 7 shows these $(2,1)$ decompositions as blue lines on the Bloch sphere surface.
Figure 3: Convexification procedure of the two corresponding pure states lying on the line given by the vector $|Z_{1}\rangle+|Z_{2}\rangle$ in a mixed state. The curves are made convex with a line touching at $p_{c}=0.962243$, as is highlighted in the inset, were we plot the differenz in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line. The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are above both curves; they are not shown here.
In Neveling and Osterloh (2024), much effort has been given concerning the distribution of optimal decompositions. These findings hold for even more general SL-invariant tangles $\tau$ than the threetangle, detected by $\sqrt{\tau_{3}}$, and the observations here agree with it. We emphasize that the simplices of $|GHZ\rangle-|Z_{1}\rangle-|Z_{2}\rangle$ is co-planar with $|Z_{1}\rangle-|Z_{2}\rangle-|N_{1}\rangle$; so are the simplices $|Z_{1}\rangle-|N_{1}\rangle-|W\rangle$ and $|Z_{1}\rangle-|N_{3}\rangle-|W\rangle$.
Optimal $(2,1)$ decompositions in between the states $|N_{i}\rangle$ In order to get a glimpse of what is happening in between the lower tetrahedra, we briefly want to see the case in which a density matrix lies within this simplex of $(2,1)$ decompositions. For symmetry reasons, this is the case for the two pure states lying respectively on the vectors corresponding to $|Z_{1}\rangle+|Z_{2}\rangle$ or to $|Z_{3}\rangle+|W\rangle$. The two cases lead to the same result, and we will describe explicitly the first case of $|Z_{1}\rangle+|Z_{2}\rangle$. The corresponding characteristic curves can be viewed in Fig. 3. It is seen that a convexification leads here to a $p_{c}=0.962243$ which corresponds to $p_{2}(p_{0},\pi,p_{c})=0.989858$ for the corresponding pure state on the Bloch sphere surface. This leads to an angle $\theta_{1}=0.201758=11.5599^{\circ}$. As before we have to consider that the two eigenstates lie at an angle $\theta_{0}=2.0539=117.68^{\circ}$. So, the angle of the state $|M_{3}\rangle$ is $\theta=\theta_{0}+\theta_{1}=2.25566=129.24^{\circ}$. The three points corresponding to the states $|M_{i}\rangle$ for $i=1,2,3$ are marked in Fig. 7 with white dots.
For the further points in the Bloch sphere, we rely on a direct comparison of $(2,1)$ with $(1,1)$ decompositions. We therefore chose the connecting line between $|N_{3}\rangle$ and $|M_{3}\rangle$; the complete solution is obtained by applying the three-fold symmetry. It is important noticing that, at first sight, the effective tangle of $(2,1)$ and $(1,1)$ decompositions yield the same result. That this is not so is seen by looking at the difference between $(1,1)$ and $(2,1)$ decompositions, shown in Fig. 4. $\lambda$ refers to the parameter convexly connecting $|N_{3}\rangle$ and $|M_{3}\rangle$, whereas $\lambda_{2}$ is the corresponding parameter for $|W\rangle$ and $|Z_{3}\rangle$.
Figure 4: Difference of $\sqrt{\tau_{3}}$ of $(1,1)$ and $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
It is seen that the states under consideration get optimised by choosing the $(2,1)$ decomposition. We optimize the tangle with respect to $\lambda_{2}$. The pure states are located in the positions indicated in Fig. 5.
Figure 5: Positions of the entangled pure state in the $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
This leads to the lower blue curves shown in Fig. 7 which lie on a circle that is not a grand circle as the ones before; instead it has a distance to the center of the Bloch sphere of $0.0711148$. There are numerical errors to this, but they are small and only occur in the border regions, where the difference to the $(1,1)$ decompositions tends to vanish (see Fig. 6)
Figure 6: Difference of $\sqrt{\tau_{3}}$ of $(1,1)$ and $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
That the states form a circle is confirmed by calculating the derivative with respect to the angle and showing its orthogonality to the original vector (not shown here). The numerical error scales with $\Delta\phi^{2}$ as expected. The normal direction of the circle is $(0.57589,0,-0.81753)$. Using the symmetry, we get the three points of normal vectors together with a dashed line in blue. The same is done for the azimuthal grand circles.
For the remaining points in the Bloch sphere there is only a single visible state from the zero polytope left which leads to $(1,1)$ decomposition being optimal there. They are depicted in Fig. 7 by green lines and red points.
Figure 7: Optimal decompositions are obtained for the rank-two mixtures of GHZ and W state and are depicted here graphically. It consists of the zero polytope (blue shaded), 4 polytopes defined by three of the zero states together with a state $|N\rangle$ one of which is the GHZ state itself (grey shaded); all these polytopes are three dimensional. On top of this kernel structure of five polytopes are three lines along grand circles along $\phi$ direction and three lines which are on circles with a distance $0.0711148$ to the center of the Bloch sphere. Along these lines move pure states which form a two-dimensional optimal decomposition (red triangles). Blue points mark the position of the respective normal vectors on the Bloch sphere surface; blue lines the corresponding pure states in $(2,1)$ decompositions. All remaining parts of the Bloch sphere are covered by $(1,1)$ decompositions.
Conclusions – The mixture of GHZ and W state has been elaborated for all the Bloch sphere, obtaining a pattern on the surface that is characteristic for these two states in the mixture. The finding is based on the zero-state locking and that $(n_{z},n_{e})$ decompositions for $n_{e}>1$ are not optimal for the present case. We have checked this numerically but there is no general proof for that finding. For the zero-state locking it is essential to look at a certain roots of the tangle, which is the square root in the case of the threetangle. The Bloch sphere is divided into the zero polytope and four polytopes onto each faces of the zero polytope with corresponding four pure states. The faces of two neighboring such polytopes are strictly planar. There are pure states that belong to $(2,1)$ decompositions moving on three azimuthal grand circles and three circles which possess a small distance $0.0711148$ to the Bloch sphere center. Their normal vector is $(0.57589\cos\phi_{i},0,-0.81753\sin\phi_{i})$, with $\phi_{i}=i2\pi/3$. The remaining decompositions to fill the Bloch sphere are of the type $(1,1)$. This distribution of the optimal decomposition is natural and inter-connects the polytopes of various dimensions. For SL invariant tangles there is a transformation rule that applies in particular to the optimal decomposition statesViehmann et al. (2012). Applying these rules yields the solution for the convex-roof for these cases. For the general case the characteristic features of the convex roof will modify, but the structure will stay the same, as reported also in ref. Neveling and Osterloh (2024). It would be interesting to what extent the findings are universal. The paths of the pure entangled state of $(2,1)$ decompositions here moves on circles, but this does not need to be this way. Future investigation, clarifying this question, will be needed. The analysis is based on two essential pillars: the zero-state locking and that every optimal decomposition would be of the type $(n_{z},1)$. It would be crucial to either prove or disprove these conjectures in future work.
This presumably clarifies how optimal decompositions of rank two mixtures behave. It will be very intriguing to understand their behavior for higher ranks.
Acknowledgements – I acknowledge the TII and Luigi Amico for supporting this research.
Appendix A General properties of optimal decompositions for real states of rank two
For the number $n_{\rm opt}$ of pure states contained in any optimal decomposition holds ${\rm rank}[\rho]\leq n_{\rm opt}\leq({\rm rank}[\rho])^{2}$ Carathéodory (1911); Uhlmann (1998). Thus, it is sufficient to look for up to $4$ such pure states that generically form three-dimensional simplices. We want to emphasize that more than the maximal $4$ states may be in an optimal decomposition; in this case, since every sub-partition of an optimal decomposition is itself optimal, each $4$ states out of that decomposition are optimal as well and the optimal decomposition is given by the convex polytope made out of these points. The tangle will behave linearly in these optimal polytopes. Since the states are composed of real elements every tangle $\tau$ has the property that $\tau(z)\equiv\tau(z^{*})$ where $|\psi\rangle\propto|\psi_{1}\rangle+z|\psi_{2}\rangle$ is a representation that corresponds to a vector onto the Bloch-sphere with the parametrizationOsterloh (2016)
$z\mathrel{\widehat{=}}\vec{n}=\left(\begin{array}[]{c}x\\ y\\ z\end{array}\right)=\left(\begin{array}[]{c}2\sqrt{p(1-p)}\cos\phi\\ 2\sqrt{p(1-p)}\sin\phi\\ 2p-1\end{array}\right)$ (4)
where $|GHZ\rangle$ and $|W\rangle$ are the two anti-podes on the z-axis. Real superpositions lie in the $x$-$z$-plane.
We will be mainly interest in a) where the line $\overline{\vec{n}_{0}\vec{n}_{1}}$ cuts the $z$-axis, b) which are the weights of the corresponding entangled state(s), and c) where it crosses the Bloch sphere, that is the corresponding pure states. Whereas the answer to a) is given by the probability
$P=\frac{p_{0}\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}+p_{1}\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{p_{0}\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}+p_{1}\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (5)
such that the point is located in $\vec{n}_{P}=(2P-1)\vec{e}_{z}$, the answer to b) is given by
$\lambda=\frac{\mu_{1}}{\mu_{0}}=\frac{\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (6)
and for the mainly interesting case of $\phi_{1}=0$ and $p$ being the probability in the density matrix in consideration, we obtain as a solution to c)
$P_{\pm}=p_{0}+\frac{p-p_{0}}{2}\frac{(p-p_{0})(1-2p_{0})+2p_{0}(1-p_{0})\cos^{2}\phi_{0}\pm\sqrt{(p_{0}-p)^{2}+4pp_{0}(1-p_{0})(1-p)\cos^{2}\phi_{0}})}{(p-p_{0})^{2}+p_{0}(1-p_{0})\cos^{2}\phi_{0}}$ (7)
The resulting tangle is a convex combination of the tangle values of the respective pure states.
Figure 8: States within the Bloch sphere. For a given line $\overline{x_{1}x_{2}}$ all triangles $z_{0}x_{1}z_{1}$, $z_{0}x_{2}z_{2}$, and $z_{0}x_{2}^{\prime}z_{2}^{\prime}$, are similar. Ratios of their length are determined by the theorem of intersecting lines.
This strictly linear behavior inside all the simplices of an optimal decomposition indirectly tells about whether there needs to be one or more states in a decomposition for being optimal: the continuation of the respective decomposition type has to lead to a convex behavior in the tangle value; where this condition is not satisfied a (linear) convexification is needed. This holds in particular if the corresponding optimal polytope changes dimension. Here, this means to a dimensionality of at most three.
A.1 $(1,1)$-decompositions
Here, we collect the formulae obtained for $(1,1)$-decompositions of a pure state at z-coordinate $12p_{1}-$ 1, where $p_{1}$ corresponds to its probability of the two states $|\psi_{i}\rangle$; $i\in\{N,S\}$. This corresponds to two pure states in the Fig. 8. We find
$l_{1}(p_{1},p)$ $=$ $2\sqrt{(p-p_{1})^{2}+p_{1}(p_{1}-1)}$ (8)
$l_{2}(p_{1},p)$ $=$ $\frac{4p(1-p)}{l_{1}(p_{1},p)}$ (9)
$m_{2}(p_{1},p)$ $=$ $\frac{l_{1}(p_{1},p)}{l_{1}(p_{1},p)+l_{2}(p_{1},p)}$ (10)
$p_{2}(p_{1},p)$ $=$ $\frac{4p^{2}(1-p_{1})}{l_{1}(p_{1},p)^{2}}$ (11)
$=$ $(1-p_{1})\frac{p}{1-p}\cdot\frac{l_{2}(p_{1},p)}{l_{1}(p_{1},p)}$ (12)
$p(p1,p2)$ $=$ $p_{1}+\frac{(p_{2}-p_{1})\sqrt{p_{1}(1-p_{1})}}{\sqrt{p_{1}(1-p_{1})}+\sqrt{p_{2}(1-p_{2})}}$ (13)
Next we want to examine whether for a given $(1,1)$ decomposition with $p_{1}$, $p_{2}$, and $p$ it is convenient to split the pure state into two, having a $(1,2)$ decomposition. It can be seen that the ratios $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ are equal with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}_{2}}^{2}}$ (see Fig. 8). For the new $p^{\prime}_{2}$ of this new state at phase $\phi_{2}$ the calculation gives
$p^{\prime}_{2}(p_{1},p;\phi_{2})=p_{1}+(p-p_{1})\frac{2p_{1}(1-p_{1})-(p-p_{1})(2p_{1}-1)\cos^{2}\phi_{2}+\cos\phi_{2}\sqrt{4p(1-p)p_{1}(1-p_{1})+(p-p_{1})^{2}\cos^{2}\phi_{2}}}{2(p_{1}(1-p_{1})+(p-p_{1})^{2}\cos^{2}\phi_{2})}$ (14)
A.2 $(2,2)$-decompositions
In the following we will assume that two complex conjugated pairs, corresponding to a superposition of two states at $p_{i}$ and angles ±$\pm\phi_{i}$, $i\in\{1,2\}$, such that both states come to lie inside the Bloch sphere projection. We find
$m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $\frac{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (15)
$=$ $\frac{l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)}{l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)+l_{2}(p_{1},\phi_{1};p_{2},\phi_{2};p)}$ (16)
$m_{1}(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $1-m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ (17)
$p(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $p_{1}+(p_{2}-p_{1})\frac{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (18)
$=$ $\frac{p_{2}\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+p_{1}\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (19)
$=$ $p_{1}m_{1}(p_{1},\phi_{1};p_{2},\phi_{2})+p_{2}m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ (20)
$l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)$ $=$ $2\sqrt{(p_{1}-p)^{2}+p_{1}(1-p_{1})\cos^{2}\phi_{1}}$ (21)
$l_{2}(p_{1},\phi_{1};p_{2},\phi_{2};p)$ $=$ $2\sqrt{(p_{2}-p)^{2}+p_{2}(1-p_{2})\cos^{2}\phi_{2}}$ (22)
From this it is seen that the weights of the states are symmetric under exchanging $p_{i}\to 1-p_{i}$ separately for $i\in\{1,2\}$. Setting $\phi_{2}=0$, we are in $(2,1)$ decompositions and the result is
$p_{2}(p_{1},\phi_{1};p)=p_{1}+(p-p_{1})\frac{(p_{1}-p)(2p_{1}-1)+2p_{1}(1-p_{1})\cos^{2}\phi_{1}+\sqrt{(p_{1}-p)^{2}+4p(1-p)p_{1}(1-p_{1})\cos^{2}\phi_{1}}}{2((p_{1}-p)^{2}+p_{1}(1-p_{1})\cos^{2}\phi_{1})}$ (23)
Next, for the evaluation of $(2,2)$ from $(2,1)$ decompositions for given $p_{1}$, $\phi_{1}$, and $p$ the ratios of the following quantities are kept fixed (see Fig. 8): $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}}_{2}^{2}}$. These have to be substituted in $m_{2}(p_{1},\phi_{1};p^{\prime}_{2},\phi^{\prime}_{2})$ where $x^{\prime}_{2}=2\sqrt{p^{\prime}_{2}(1-p^{\prime}_{2})}\cos\phi^{\prime}_{2}$.

### v2

@@
 Abstract
-I present an exact solution for the convex roof of the square root threetangle for all states within the Bloch sphere.
-The working hypothesis is that optimal decompositions contain as many states from the zero-polytope as possible which can be called zero-state locking.
+I present an exact solution for the convex roof of the square root of the threetangle for all states within the Bloch sphere.
+The working horse that optimal decompositions contain as many states from the zero-polytope as possible which can be called zero-state locking is proved and an inequality is derived which decides about the optimality of the decompositions under consideration here.
 The footprint of the measure of entanglement consists in a characteristic pattern for the fixed pure states on the surface which form the optimal solution.
@@
 I Introduction
-Although we know that entanglement is everywhere present where there is interaction, it is difficult to measure it, because the operator one has to measure is non-Hermitean and consequently the eigenvalues are complex values for every pure states.
-We therefore take its absolute value to quantify this important resource, which is indispensable for quantum computingLadd et al.
+Entanglement is omnipresent in the presence of interaction and it is difficult to measure, because the operator to measure is non-Hermitean, a case rarely encountered in physical systems.
+Its expectation value is not calculated as the trace with the density matrix of the state in consideration; instead it is complicated by the convex-roofVidal (2000) which in general is NP-hard to calculateGurvits (2003, 2004).
+It is an important quantity, indispensable for quantum computingLadd et al.
 (2010); Gill et al.
@@
 (2018); Polino et al.
-(2020); quantum technology is working because these non-local quantum correlations are there.
-Reasonably, as a non-local quantity, entanglement has to be invariant under the change of the local basis and permutations of the local entities.
+(2020); quantum technology works through these nonlocal quantum correlations.
+As a non-local quantity, entanglement has to be invariant under the change of the local basis and permutations of the local entities.
 This leads to the concept of LOCC invariance and the monotone propertyVidal (2000) that every good entanglement measure must satisfy.
-In general one has to prove this monotone property, which is granted for, when extending the local invariance group from $SU(n)$ to $SL(n)$Verstraete et al.
+In general, this monotone property needs a proof, which however is granted for by extending the local invariance group from $SU(n)$ to SL(n)SL(n)Verstraete et al.
 (2003) for $n$ being the number of different axis of the local entity one can do rotations about.
-In general, the quantities obtained through this extension have the elegant property of being merely susceptible for entanglement at a global level, discarding everything which is non-global.
-Thus, they are measures of genuine multipartite SL-entanglementOsterloh and Siewert (2005); Doković and Osterloh (2009).
-Examples for taking qubits as local entities are the concurrenceHill and Wootters (1997); Wootters (1998) and the threetangleCoffman et al.
+The quantities obtained through this extension have the elegant property of being essentially susceptible to entanglement at a global level, discarding everything which is non-global.
+They are measures of genuine multipartite SL-entanglementOsterloh and Siewert (2005); D oković and Osterloh (2009); Neveling and Osterloh (2024).
+Simple examples for qubits as local entities are the concurrenceHill and Wootters (1997); Wootters (1998) and the threetangleCoffman et al.
 (2000).
-The monotone property has included already the convex-roof extension to arbitrary mixed states.
-However, the calculation of it turns out to be a difficult task, which even is thought to be NP-hardGurvits (2003, 2004) (see however Wen et al.
+Figure 1: Optimal decompositions are obtained for the rank-two mixtures of GHZ and W state and are depicted here graphically.
+It consists of the zero polytope (blue shaded), 4 polytopes defined by three of the zero states together with a state $|N\rangle$ one of which is the GHZ state itself (grey shaded); all these polytopes are three dimensional.
+On top of this kernel structure of five polytopes are three lines along grand circles along $\phi$ direction and three lines which are on circles with a distance $0.0711148$ to the center of the Bloch sphere.
+Along these lines move pure states which form a two-dimensional optimal decomposition (red triangles).
+Blue points mark the position of the respective normal vectors on the Bloch sphere surface; blue lines the corresponding pure states in $(2,1)$ decompositions.
+All remaining parts of the Bloch sphere are covered by $(1,1)$ decompositions.
+Included in the monotone property is the convex-roof extension to arbitrary mixed states, which can be NP-hardGurvits (2003, 2004).
+For very few examples we have an exact solution for evaluating this convex-roof, as for the concurrence for two qubits.
+These are examples for which the pure state measure scales linear in $\rho$; it generally holds if the corresponding $d$-th root of a tangle of degree $2d$ is connected to a bilinear functional as is the concurrence.
+This happens e.g.
+for the threetangle for certain statesNeveling and Osterloh (2024) (see also the telescope states in Eltschka et al.
 (2009)).
-For very few examples we have an exact solution for evaluating this convex-roof, as for the concurrence for two qubits.
-These are examples for which the pure state measure scales linear in $\rho$; it generally holds if the corresponding $d$-th root of a tangle of degree $2d$ is connected to a bilinear functional as does the concurrence.
-This happens for the threetangle but only for certain statesNeveling and Osterloh (2024).
-This is not so in general, and the zero-polytope consists of four independent solutions.
-However, for the mixture of GHZ and W states and the square root of the threetangle it is knownLohmayer et al.
+One exclusive property of SL-invariant entanglement measures is that they are given as multi-nomials of complex variables $z_{i}$.
+The superpositions of two states having zero tangle of degree $2d$ is given by the $2d$ solutions of a polynomial of degree $2d$.
+These zero-states form a polytope with $2d$ corners in the Bloch sphere picture: the zero-polytope.
+In general the zero-polytope consists of $2d$ independent solutions.
+It is knownLohmayer et al.
 (2006); Eltschka et al.
 (2008); Viehmann et al.
-(2012) that the Bloch-sphere has two tetrahedra in it: one is the zero-polytope and the other is the polytope made of three of the states of the zero-polytope and the GHZ state.
-Based on the conjectured property of zero-state lockingOsterloh (2016); Neveling and Osterloh (2024), in a recent workNeveling and Osterloh (2024) it has been established how optimal decompositions will behave for rank-$2$ density matrices (see Fig.
-1).
-Optimal decompositions are limited to at most $4$ statesCarathéodory (1911); Uhlmann (1998), i.e.
-to three dimensional simplices for this case.
-Figure 1: Polytopes of type $3^{+}N$ (left figure) and $3^{+}S$ (right figure), taken from Ref.
-Neveling and Osterloh (2024).
-To the zero polytope (blue) entangled polytopes (gray) are added with corresponding pure states $|N_{\pm}\rangle$.
-It was noted that a pure state will be added for every of the faces of the zero polytope, which are four in this particular case.
-For symmetry requirements $|N_{\pm}\rangle$, as well as the pure states of the $(2,1)$ decompositions, are situated in $\phi=0$.
-This condition will be relaxed with the symmetry here, because of the non-intersection property of optimal decompositions.
-The optimal decompostions, here indicated by green lines, brachiate through this structure of polytopes with the additional pure state at $\phi=0$.
-We use the nomenclature $(n_{z},n_{e})$ of this work for identifying the decompositions, where $n_{z}$ is the number of states from the zero polytope, and $n_{e}$ that of entangled pure states in the decomposition.
+(2012) for the mixture of GHZ and W states and the square root of the threetangle that the Bloch-sphere contains two tetrahedra: one is the zero-polytope and the other is the polytope made of three of the states of the zero-polytope and the GHZ state.
+In a recent workNeveling and Osterloh (2024) it has been established how optimal decompositions will behave for rank-$2$ density matrices.
+A proof of the zero-state lockingOsterloh (2016); Neveling and Osterloh (2024) crucial also for these result, is presented in the next section.
+Optimal decompositions do not intersect and are here limited to at most $4$ statesCarathéodory (1911); Uhlmann (1998) (see Appendix B), i.e.
+to three dimensional simplices.
+We use the nomenclature $(n_{z},n_{e})$ from Ref.
+Neveling and Osterloh (2024) for identifying the decompositions: $n_{z}$ is the number of states from the zero-polytope, and $n_{e}$ that of entangled pure states in the decomposition.
 In particular, it has been observed that for each three surface elements of the zero polytope there is precisely one entangled pure state $|N\rangle$ forming a $(3,1)$ polytope.
@@
 These pure states $|N_{i}\rangle$ need to be found as a convexification point of $(1,1)$ and $(2,1)$ decompositions.
-Optimal decompositions do not intersect, and this fills out good part of the Bloch-sphere.
-Then, for the tips of each two connected surface elements, $|N_{i_{1}}\rangle$ and $|N_{i_{2}}\rangle$, out of the pure entangled states $|N_{i}\rangle$ found before, are connected by a line of pure entangled states.
-These give further $(2,1)$ decompositions in the Bloch sphere.
+This fills out good part of the Bloch-sphere.
+The tips of each two states $|N_{i_{1}}\rangle$ and $|N_{i_{2}}\rangle$ are connected by a line of pure entangled states.
+They define further $(2,1)$ decompositions in the Bloch sphere.
 All remaining decompositions will be of the type $(1,1)$.
-This excludes the possibility of optimal $(0,n_{e})$ decompositions for $n_{e}>1$; there are strong arguments for this to holdNeveling and Osterloh (2024), but no general proof exists.
-Optimality of $(n_{z},n_{e})$ have been excluded numerically through case studiesNeveling and Osterloh (2024).
-The transformation property, also of optimal decomposition statesViehmann et al.
-(2011, 2012), is an important finding and lifts the results of this work to all states being SL-equivalent to the mixed state in consideration.
+This excludes the possibility of optimal $(0,n_{e})$ decompositions for $n_{e}>1$ here.
+Optimality of $(n_{z},n_{e})$ can be excluded numerically through case studiesNeveling and Osterloh (2024).
+Here, an inequality is given in Eq.
+6 where this behavior changes.
+The $SL$-invariance of the tangle induces transformation properties, also of optimal decomposition statesViehmann et al.
+(2011, 2012) and lifts the results of this work to all states being SL-equivalent to the mixed state in consideration.
 This work is laid out as follows.
-At first we will deduce those pure states $|N_{i}\rangle$ which form optimal $(3,1)$ decompositions, then find the lines interconnecting each pair of the entangled pure states $|N_{i}\rangle$.
+We will start with the presentation of the proof of the zero-state locking theorem underlying this work.
+Next, we analyze the specific $(2,1)$ and $(0,2)$ decompositions and obtain an inequality that decides about which is optimal.
+In the last section we go through the steps needed in general and for this particular case of GHZWGHZ-W mixtures of three qubits and its threetangle.
+The calculations necessary for each step is documented in the appendix.
+The result are $|N_{i}\rangle$ which form optimal $(3,1)$ decompositions and lines interconnecting each pair of the entangled pure states $|N_{i}\rangle$.
 The remaining space of the Bloch-sphere is filled with $(1,1)$ decompositions.
-The pure entangled states $|N_{i}\rangle$ We analyze the Bloch sphere of the rank-two mixed state
+Zero-state locking – We give a proof of the zero-state locking behavior for optimal decompositions.
+It states that every zero-state visible from a density matrix $\rho$ is within the optimal decomposition.
+Let $\tau_{d}$ be an SL-invariant measure of entanglement of polynomial degree $2d$; we have $\tau_{d}[\alpha\psi]=\alpha^{2d}\tau_{d}[\psi]$.
+Hence its $d$-th root scales quadratic in the wave function coefficients as it does for the concurrence.
+Thus, we take $\sqrt[d]{\tau_{d}}$ as a proper measure because of its linear scaling behavior in the probabilities Viehmann et al.
+(2012).
+The equation $\tau_{d}[\psi]=0$ has for SL-invariant tangles $\tau_{d}$ exactly $2d$ solutions.
+Each single root $z_{j}$ leads to a scaling behavior $(z-z_{j})^{1/d}$.
+Hence multiple roots scale as $(z-z_{j})^{m/d}$ if $m$ is the multiplicity [^1] I emphasize that $z$ is the stereographic projection of the Bloch sphere.
+In the $z_{j}$ the tangle $\tau_{d}$ is represented by $\tau_{d}=\prod_{j=1}^{2d}(z-z_{j})$.
+.
+Let for the locking property be $m<d$.
+We therefore have a strictly concave behavior around $z_{j}$.
+We next consider a decomposition in which some of the decomposition states are of the solutions to the zero-polytope.
+Its average tangle is $T(x_{0})=w(x_{0})e(x_{0})=w_{0}e_{0}$ where $w(x)$ is the weight of the entangled density matrix and $e(x)$ its average entanglement.
+Both the weight and the entanglement are locally holomorphic functions which have a Taylor expansion.
+Now assume that one solution of the zero-polytope be shifted by a small amount $\vec{\varepsilon}=(\varepsilon_{1},\varepsilon_{2})$ on the Bloch sphere surface [^2] The two variables could for example be $\varphi$ and $\theta$.
+The stereographic projection would lead to a stretching of $\theta$ by the value $1+|z_{j}|^{2}/4$.
+We therefore prefer to leave it more general..
+The remaining states stay unaltered.
+This consequently leads to a change of both the weight function $w(\vec{x}_{0}+\vec{\varepsilon})=w_{0}-\vec{w}_{1}\cdot\vec{\varepsilon}+\varepsilon_{i}w_{2;ij}\varepsilon_{j}+\dots$ and the entanglement function $e(\vec{x}_{0}+\vec{\varepsilon})=e_{0}+\vec{e}_{1}\cdot\vec{\varepsilon}+\varepsilon_{i}e_{2;ij}\varepsilon_{j}+\dots$ such that the average entanglement of the new decomposition results in
+$T(\vec{x}_{0}+\vec{\varepsilon})$ $=$ $(1-w(\vec{x}_{0}+\vec{\varepsilon}))e_{z}|\vec{\varepsilon}|^{m/d}$
+$\ +w(\vec{x}_{0}+\vec{\varepsilon})e(\vec{x}_{0}+\vec{\varepsilon})$
+$\approx$ $w_{0}e_{0}+(1-w_{0})e_{z}|\vec{\varepsilon}|^{m/d}$
+$\quad+(w_{0}\vec{e}_{1}-e_{0}\vec{w}_{1})\cdot\vec{\varepsilon}$
+with $e_{z}$ a positive prefactor.
+For $m<d$ in the limit $\varepsilon\to 0$ one finds $T(\vec{x}_{0}+\vec{\varepsilon})=w_{0}e_{0}+(1-w_{0})e_{z}|\vec{\varepsilon}|^{m/d}$ which is larger than $T(x_{0})$ as long as the condition $m<d$ is satisfied.
+This property excludes the possibility of optimal $(0,n_{e})$ decompositions if some of the zero-states are visible from the density matrix in consideration (optimal decomposition polytopes can be viewed as intransparent for density matrices, since optimal decomposition cannot intersect).
+For $m=d$ one obtains $(w_{0}\vec{e}_{1}-e_{0}\vec{w}_{1})\cdot\vec{\varepsilon}+(1-w_{0})e_{z}|\vec{\varepsilon}|$.
+This is still positive and the optimal decomposition would stay zero-locked unless $(w_{0}\vec{e}_{1}-e_{0}\vec{w}_{1})\cdot\vec{\varepsilon}<-(1-w_{0})e_{z}|\vec{\varepsilon}|$.
+Curiously, the integrable case where we have $\sqrt[d]{\tau_{d}[z]}=(z-z_{1})(z-z_{2})$, hence both zeros in the system have a multiplicity of $d$.
+Here, however, the inequality has to become an equality because no zero-state locking is observed and the optimal decomposition can be chosen arbitrarily as in the case of the concurrence (see also Ref.
+Uhlmann (2000)) where decompositions can be made of equally entangled pure statesHill and Wootters (1997); Wootters (1998).
+Whenever $m>d$, there is no zero-state locking for this particular state from the zero-polytope.
+For the rest it nevertheless applies.
+Because optimal decompositions must be continuous and no two optimal decompositions can intersect each other (unless the pure states of the decompositions are joined together) this zero-state locking behavior excludes optimal decompositions without visible states from the zero-polytope.
+$(n_{z},n_{e})$-decompositions with $n_{e}>1$ – Even if the zero-state locking applies to $z$, it is not clear that there are no $(0,n_{e})$ decompositions to fill up the Bloch sphere.
+Assume there would exist $(n_{z},n_{e})$-decompositions.
+Its shadow would leave a dark space in the Bloch sphere.
+Optimal decompositions cannot intersect.
+This means that every simplex can be viewed as intransparent, leaving the states that can be viewed from a density matrix as the only possible pure states for an optimal decomposition.
+In this specific case we can limit ourselves with $(n_{z},2)$-decompositions with $1\leq n_{z}\leq 2$ being visible and subject to zero-state locking.
+If $n_{z}=1$ there are states in the Bloch sphere where the zero-state cannot be seen.
+For these states one would necessarily obtain optimal decompositions in which no zero-state participates.
+This would however not contradict the zero-state locking because no zero-state is visible.
+For $n_{z}=2$ we obtain an analoguous statement.
+Summarizing, if $(n_{z},n_{e})$ decompositions exist with $n_{e}>1$ so there need to be $(0,n_{e})$ decompositions in the Bloch sphere.
+Figure 2: Left: The zero-polytope is shown (blue) in the Bloch sphere projection in $(x,z)$ plane, together with the sphere of radius $r_{0}$ given by the two complex conjugated yero states.
+The entangled pure states marking optimal decompositions are marked by green dots.
+The plane inclined by an angle $\delta_{1}$ is shown in purple color, cutting both spheres in circles of radii $d_{0}$ and $d=d_{0}+d_{1}$.
+The inclination angle is limited by the three-dimensional polytopes by the interval $\delta_{-},\delta_{+}+\pi/2]$.
+$\delta_{-}$ is negative.
+Right: The purple plane inclined by the angle $\delta$ in the Bloch sphere is shown with part of the circles of radii $d_{0}$ and $d=d_{0}+d_{1}$.
+The density matrix $\rho$ decomposed of the two complex conjugated states located in $(x,\pm y)$ must be better than the $(2,1)$ decomposition in consideration.
+For any $(2,2)$ decomposition being optimal for a state $\rho^{\prime}$ located at distance drdr from the point $z_{0}$, there need to be states $\rho$ that have to be optimally decomposed by two corresponding states in $(x,\pm y)$.
+Assume that we have a situation of real wavefunction coefficients, as one has in the unique ground state of a Hamilonian.
+This implies a further symmetry: the solutions to the zero polytope are constituted of complex conjugated pairs.
+Let two such solutions be $z$ and $z^{*}$ with a real part $z_{0}=(x_{0},0,y_{0})$ which has a distance $r_{0}=\sqrt{x_{0}^{2}+y_{0}^{2}}$ to the origin (see left panel of Fig.
+2.
+The azimuthal angle is $\vartheta=\arctan(x_{0}/y_{0})$.
+The normalized vector $\vec{n}=(\cos(\vartheta_{1}),\sin(\vartheta_{1}))$ describes the direction of the projected plane (purple line/plane ) onto what corresponds to the real plane in Hilbert space.
+This line $\vec{z}_{0}+d~\vec{n}$ hits the Bloch sphere surface in the two points
+$d_{1/2}=-r_{0}|\cos(\vartheta-\vartheta_{1})|\pm\sqrt{1-r_{0}^{2}\sin^{2}(\vartheta-\vartheta_{1})}\;$ (3)
+where $\rho_{1}=r_{0}|\cos(\vartheta-\vartheta_{1})|$ is the radius of the inner and $\rho=\sqrt{1-r_{0}^{2}\sin^{2}(\vartheta-\vartheta_{1})}$ of the outer circle (see right panel of Fig.
+2).
+Defining $\Delta x(\delta\varphi)$ such that $x(\delta\varphi)+\Delta x(\delta\varphi)=d_{1}$, a straight forward calculation yields
+$\left(\begin{array}[]{c}\Delta x(\delta\varphi)\\ y(\delta\varphi)\end{array}\right)=\rho\left(\begin{array}[]{c}1-\cos\,\delta\varphi\\ \sin\,\delta\varphi\end{array}\right)\approx\rho\,\left(\begin{array}[]{c}\delta\varphi^{2}/2\\ \delta\varphi\end{array}\right)$ (4)
+In this plane the tangle will behave as
+$\tau\approx\tau_{0}+\frac{\tau"(0)}{2}\delta\varphi^{2}$ (5)
+If $\tau"(0)>0$ there is a minimum of $\tau$ and the $(2,1)$ decomposition is clearly better.
+Hence we assume that $\tau"(0)<0$.
+For the $(2,1)$ decomposition we obtain $\tau_{(2,1)}=\tau_{0}(d_{1}-\Delta x(\delta\varphi))/d_{1}$ and for the $(0,2)$ decomposition $\tau_{(0,2)}=\tau_{0}(1-\frac{|\tau"(0)|}{2\tau_{0}}\delta\varphi^{2})$.
+Thus $\tau_{(2,1)}-\tau_{(0,2)}\approx(|\tau"(0)|-\tau_{0}\frac{\rho}{d_{1}})$.
+This difference vanishes iff $|\tau"(0)|=\tau_{0}\frac{\rho}{d_{1}}$.
+It is positive iff $\tau_{(0,2)}<\tau_{(2,1)}$.
+This leads to
+$|\tau"(0)|>\tau_{0}\frac{\rho}{d_{1}}\;.$ (6)
+Figure 3: Difference $\Delta\tau:=\tau_{(0,2)}-\tau_{(2,1)}$ of the (0,2) and $(2,1)$ decompositions is shown for $\tau:=\sqrt{\tau_{3}}$ is shown depending on $\delta_{1}$ for $p$ close to $1$.
+Here, the tangle is the square-root of the threetangle.
+With $|GHZ\rangle=(|000\rangle+|111\rangle)/\sqrt{2}$ and $|W\rangle=(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$ being the poles of the Bloch sphere, the pure states’ tangle is
+$\sqrt{|\tau_{3}|}(p,\varphi)$ $=$ $\left|p^{2}+\frac{2^{7/2}}{3^{3/2}}\sqrt{p(1-p)^{3}}e^{i3\varphi}\right|^{1/2}$ (7)
+$=$ $\left[p^{4}+\frac{32p^{2}}{\sqrt{54}}\sqrt{p(1-p)^{3}}\cos 3\varphi\right$. (8)
+$\qquad\qquad\qquad\quad+\left.\frac{16^{2}}{54}p(1-p)^{3}\right]^{1/4}$
+In the characteristic curves in Ref.
+Lohmayer et al.
+(2006) (taking the square root) it is seen that the main $\varphi$ dependence originates at $p_{0}$.
+In Fig.
+3 the behavior is shown for all admissible $\delta_{1}$.
+We obtain
+$\sqrt{|\tau_{3}|}(p_{0},\delta\varphi)=\sqrt{2}p_{0}\left(1-\frac{3^{2}}{4^{2}}\delta\varphi^{2}\right)\;.$ (9)
+We have $|\tau"(p_{0},0)|=\frac{9}{8}\tau(p_{0},0)<2\tau(p_{0},0)$ since $\rho=2d_{1}$ for this case. Hence, the $(2,1)$ decomposition is optimal.
+General procedure – In Neveling and Osterloh (2024), much effort has been put into the distribution of optimal decompositions.
+These findings hold for even more general SL-invariant tangles $\tau$.
+We describe here the general procedure for an SL-invariant n-tangle of degree $2d$ given a rank two density matrix $\rho$ with eigenstates $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$.
+At first the zero-polytope has to be calculated in solving $\tau[|\psi_{0}\rangle+z|\psi_{1}\rangle]=P_{2d}[z]=0$ with all states whose n-tangle is zero as corner states.
+Here, $P_{2d}$ is a polynomial of degree $2d$.
+The n-tangle is zero within this convex polytope and non-zero outside.
+The polytope has two dimensional simplices as faces, which comprises three pure states; for each face one has to find the entangled pure state of the optimal decomposition states and the lines interconnecting them.
+This structure of points and lines is the signature of the optimal decomposition states of the specific n-tangle to all rank-two density matrices composed of the states $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$.
+We analyze the Bloch sphere of the rank-two mixed state
+$\rho[p]=p|GHZ\rangle\langle GHZ|+(1-p)|W\rangle\langle W|$ (10)
+where $|GHZ\rangle=(|111\rangle+|000\rangle)/\sqrt{2}$ and $|W\rangle=(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$.
+A three-fold rotational symmetry due to cyclic permutations of the qubits exist for both states.
+Therefore, it is enough to perform the classification in a wedge of the Bloch sphere given by the azimuthal angle $\phi\in[-\pi/3,\pi/3]$ only.
+The rest of the sphere is obtained by symmetry requirements.
+Superposition can be given as $|\Psi[z]\rangle:=|W\rangle+z|GHZ\rangle$ with a complex number $z$.
+The zero-polytope is obtained by solving $\tau_{3}[z]:=\tau_{3}[\Psi[z]]=0$.
+All wave function coefficients are real and consequently we have $\tau_{3}[\Psi[z]]=\tau_{3}[\Psi[z^{*}]]^{*}$.
+Therefore we have a further symmetry, namely $|\tau_{3}|[z]=|\tau_{3}|[z^{*}]$.
+This facilitates the search for the structure of the convex-roof since both the state $N_{1}$ and the line intersecting $|N_{1}\rangle$ with $|GHZ\rangle$ must lie in the real plane.
+The pure entangled states $N_{i}$ – One of the pure entangled states is the maximally entangled GHZ stateLohmayer et al.
+(2006); Viehmann et al.
+(2012).
+This is also confirmed through the convexification procedure of $(2,1)$ and $(1,1)$ decompositions which correctly points to the GHZ state.
+Remains to find the state in one of the three vertical wedges of the Bloch sphere.
+It is worth choosing another orthogonal basis for this purpose.
+The orthogonal decomposition chosen here is between the real one of the three pure states of the zero-polytope and its orthogonal partner.
+These are given by
+$|Z_{3}\rangle$ $=$ $\sqrt{p_{0}}|GHZ\rangle-\sqrt{1-p_{0}}|W\rangle$ (11)
+$|Z_{3;\perp}\rangle$ $=$ $\sqrt{1-p_{0}}|GHZ\rangle+\sqrt{p_{0}}|W\rangle$ (12)
+with $p_{0}=\frac{4\sqrt[3]{2}}{3+4\sqrt[3]{2}}$ Lohmayer et al.
+(2006) (the calculations are shown in B).
+The angle of the new state is $\theta=1.99158=114.109^{\circ}$.
+The four tetrahedra are shown in grey in Fig.
+1 with green edges.
+Invoking the three-fold symmetry for this case leads to the three optimal $(3,1)$ decompositions with the states $|N_{i}\rangle$, $i=1,2,3$, as a tip.
+Consideration of $(n_{z},2)$ decompositions of the previous section demonstrates that in this case $(2,1)$ decompositions are optimal in the direction of $|GHZ\rangle$ and $(1,1)$ decompositions are optimal in the direction of $|W\rangle$.
+Fig 1 shows these $(2,1)$ decompositions as blue lines on the Bloch sphere surface.
+We emphasize that the simplices of $|GHZ\rangle-|Z_{1}\rangle-|Z_{2}\rangle$ is co-planar with $|Z_{1}\rangle-|Z_{2}\rangle-|N_{1}\rangle$; so are the simplices $|Z_{1}\rangle-|N_{1}\rangle-|W\rangle$ and $|Z_{1}\rangle-|N_{3}\rangle-|W\rangle$.
+Optimal $(2,1)$ decompositions in between the states $|N_{i}\rangle$ In order to get a glimpse of what is happening in between the lower tetrahedra, we briefly want to consider the case in which an orthogonal decomposition of the density matrix lies within this simplex of $(2,1)$ decompositions.
+For symmetry reasons, this is the case for the two pure states lying respectively on the vectors corresponding to $|Z_{1}\rangle+|Z_{2}\rangle$ or to $|Z_{3}\rangle+|W\rangle$.
+The two cases lead to the same result (for the calculations see C).
+The angle of the state $|M_{3}\rangle$ is $\theta=2.25566=129.24^{\circ}$.
+The three points corresponding to the states $|M_{i}\rangle$ for $i=1,2,3$ are marked in Fig.
+1 with white dots.
+For the lines on the Bloch sphere, we rely on a direct comparison of $(2,1)$ with $(1,1)$ decompositions.
+We therefore chose the connecting line between $|N_{3}\rangle$ and $|M_{3}\rangle$; the complete solution is obtained by applying the symmetries and the details are described in C.
+This leads to the lower blue curves shown in Fig.
+1 which lie on a circle with distance to the center of the Bloch sphere being $0.0711148$.
+That the states form a circle is confirmed by calculating the derivative with respect to the angle and showing its orthogonality to the original vector (not shown there).
+The numerical error scales with $\Delta\phi^{2}$ as expected.
+The normal direction of the circle is (0.57589,0,0.81753)(0.57589,0,-0.81753).
+Using the symmetry, we get the three normal vectors given by dashed line in blue to the blue points on the sphere and for the azimuthal grand circles as well.
+For the remaining parts of the Bloch sphere there is only a single visible state from the zero polytope left and leads to $(1,1)$ decomposition being optimal.
+They are depicted in Fig.
+1 by green lines and red points.
+Conclusions – The mixture of GHZ and W state has been elaborated for all the Bloch sphere, obtaining a pattern on the surface that is characteristic for these two states in the mixture.
+We give a proof of the zero-state locking where the finding is based on.
+We also derive an inequality which decides whether $(n_{z},n_{e})$ decompositions for $n_{e}>1$ are optimal and find them not optimal for the present case.
+For the zero-state locking it is essential to look at a certain roots which render the tangle to scale linear with the density matrix.
+This is the square root in case of the threetangle.
+The Bloch sphere is divided into the zero polytope and four polytopes onto each faces of the zero polytope with corresponding four pure states.
+The faces of two neighboring such polytopes are strictly planar.
+There are pure states that belong to $(2,1)$ decompositions moving on three azimuthal grand circles and three circles which possess a small distance $0.0711148$ to the Bloch sphere center.
+Their normal vector is $(0.57589\cos\phi_{i},0,-0.81753\sin\phi_{i})$, with $\phi_{i}=i2\pi/3$.
+The remaining decompositions to fill the Bloch sphere are of the type $(1,1)$.
+This distribution of the optimal decomposition inter-connects the polytopes of various dimensions.
+For SL invariant tangles there is a transformation rule that applies in particular to the optimal decomposition statesViehmann et al.
+(2012).
+Applying these rules yields the solution for the convex-roof for the transformed density matrix.
+For the general case the characteristic features of the convex roof will modify, but the structure will stay the same, as reported also in ref.
+Neveling and Osterloh (2024).
+It would be interesting to what extent the findings are universal.
+The paths of the pure entangled state of $(2,1)$ decompositions here moves on circles, but this does not need to be general.
+It is intriguing to look for examples where the inequality points towards $(0,2)$ decompositions in parts of the Bloch sphere.
+Future investigations, clarifying these questions, will be needed.
+This work enters the way how optimal decompositions of rank two mixtures behave.
+It will be a challenge to understand their behavior for higher ranks.
+This can also become relevant for various quantum technologies.
+Future work should establish the specific role of genuine SL-invariant entanglement in this respect.
+Acknowledgements – I acknowledge the TII and Luigi Amico for supporting this research.
+Appendix A Detecting genuine three-partite entanglement via the threetangle
+We will consider $\sqrt{|\tau_{3}|}$ as entanglement measure, where the threetangle $|\tau_{3}|$ has been defined asCoffman et al.
+(2000) (see also in Refs.
+Wong and Christensen (2001); Verstraete et al.
+(2003); Osterloh and Siewert (2005))
+$\tau_{3}$ $=$ $d_{1}-2d_{2}+4d_{3}$
+$d_{1}$ $=$ $\psi^{2}_{000}\psi^{2}_{111}+\psi^{2}_{001}\psi^{2}_{110}+\psi^{2}_{010}\psi^{2}_{101}+\psi^{2}_{100}\psi^{2}_{011}$
+$d_{2}$ $=$ $\psi_{000}\psi_{111}\psi_{011}\psi_{100}+\psi_{000}\psi_{111}\psi_{101}\psi_{010}$
+$+\psi_{000}\psi_{111}\psi_{110}\psi_{001}+\psi_{011}\psi_{100}\psi_{101}\psi_{010}$
+$+\psi_{011}\psi_{100}\psi_{110}\psi_{001}+\psi_{101}\psi_{010}\psi_{110}\psi_{001}$
+$d_{3}$ $=$ $\psi_{000}\psi_{110}\psi_{101}\psi_{011}+\psi_{111}\psi_{001}\psi_{010}\psi_{100}\ \ ,$
+and coincides with the three-qubit hyperdeterminantCayley (1846); Miyake and Wadati (2002).
+It detects states from the only genuine SL-entangled GHZ-class.
+$W$-states are not detected; they are instead detected as entangled by the pairwise Concurrence which is distributed along all possible pairs in the state.
+It is therefore not bipartite and is called genuinely multipartite entangled over the three qubits.
+However, it is only genuinely pairwise SL-entangled.
+Appendix B General properties of optimal decompositions for real states of rank two
+For the number $n_{\rm opt}$ of pure states contained in any optimal decomposition holds ${\rm rank}[\rho]\leq n_{\rm opt}\leq({\rm rank}[\rho])^{2}$ Carathéodory (1911); Uhlmann (1998).
+Thus, it is sufficient to look for up to $4$ such pure states that generically form three-dimensional simplices.
+We want to emphasize that more than the maximal $4$ states may be in an optimal decomposition; in this case, since every sub-partition of an optimal decomposition is itself optimal, each $4$ states out of that decomposition are optimal as well and the optimal decomposition is given by the convex polytope made out of these points.
+The tangle will behave linearly in these optimal polytopes.
+Since the states are composed of real elements every tangle $\tau$ has the property that $\tau(z)\equiv\tau(z^{*})$ where $|\psi\rangle\propto|\psi_{1}\rangle+z|\psi_{2}\rangle$ is a representation that corresponds to a vector onto the Bloch-sphere with the parametrizationOsterloh (2016)
+$z\mathrel{\widehat{=}}\vec{n}=\left(\begin{array}[]{c}x\\ y\\ z\end{array}\right)=\left(\begin{array}[]{c}2\sqrt{p(1-p)}\cos\phi\\ 2\sqrt{p(1-p)}\sin\phi\\ 2p-1\end{array}\right)$ (13)
+where $|GHZ\rangle$ and $|W\rangle$ are the two anti-podes on the z-axis. Real superpositions lie in the $x$-$z$-plane.
+We will be mainly interest in a) where the line $\overline{\vec{n}_{0}\vec{n}_{1}}$ cuts the $z$-axis, b) which are the weights of the corresponding entangled state(s), and c) where it crosses the Bloch sphere, that is the corresponding pure states.
+Whereas the answer to a) is given by the probability
+$P=\frac{p_{0}\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}+p_{1}\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{p_{0}\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}+p_{1}\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (14)
+such that the point is located in $\vec{n}_{P}=(2P-1)\vec{e}_{z}$, the answer to b) is given by
+$\lambda=\frac{m_{1}}{m_{0}}=\frac{\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (15)
+where $m_{i}$ is the weight of the corresponding state.
+For the mainly interesting case of $\phi_{1}=0$ and $p$ being the probability in the density matrix in consideration, we obtain as a solution to c)
+$P_{\pm}=p_{0}+\frac{p-p_{0}}{2}\frac{(p-p_{0})(1-2p_{0})+2p_{0}(1-p_{0})\cos^{2}\phi_{0}\pm\sqrt{(p_{0}-p)^{2}+4pp_{0}(1-p_{0})(1-p)\cos^{2}\phi_{0}})}{(p-p_{0})^{2}+p_{0}(1-p_{0})\cos^{2}\phi_{0}}$ (16)
+The resulting tangle is a convex combination of the tangle values of the respective pure states.
+Figure 4: Left: States within the Bloch sphere.
+For a given line $\overline{x_{1}x_{2}}$ all triangles $z_{0}x_{1}z_{1}$, $z_{0}x_{2}z_{2}$, and $z_{0}x_{2}^{\prime}z_{2}^{\prime}$, are similar.
+Ratios of their length are determined by the theorem of intersecting lines.
+Right: The optimal decompositions in the Bloch sphere spanned by GHZGHZ and $W$ state.
+This strictly linear behavior inside all the simplices of an optimal decomposition indirectly tells about whether there needs to be one or more states in a decomposition for being optimal: the continuation of the respective decomposition type has to lead to a convex behavior in the tangle value; where this condition is not satisfied a (linear) convexification is needed.
+This holds in particular if the corresponding optimal polytope changes dimension.
+Here, this means to a dimensionality of at most three.
+B.1 $(1,1)$-decompositions
+Here, we collect the formulae obtained for $(1,1)$-decompositions of a pure state at z-coordinate $2p_{1}-1$, where $p_{1}$ corresponds to its probability of the two states $|\psi_{i}\rangle$; $i\in\{N,S\}$.
+This corresponds to two pure states in the left panel of Fig.
+4.
+We find
+$l_{1}(p_{1},p)$ $=$ $2\sqrt{(p-p_{1})^{2}+p_{1}(p_{1}-1)}$ (17)
+$l_{2}(p_{1},p)$ $=$ $\frac{4p(1-p)}{l_{1}(p_{1},p)}$ (18)
+$m_{2}(p_{1},p)$ $=$ $\frac{l_{1}(p_{1},p)}{l_{1}(p_{1},p)+l_{2}(p_{1},p)}$ (19)
+$p_{2}(p_{1},p)$ $=$ $\frac{4p^{2}(1-p_{1})}{l_{1}(p_{1},p)^{2}}$ (20)
+$=$ $(1-p_{1})\frac{p}{1-p}\cdot\frac{l_{2}(p_{1},p)}{l_{1}(p_{1},p)}$ (21)
+$p(p1,p2)$ $=$ $p_{1}+\frac{(p_{2}-p_{1})\sqrt{p_{1}(1-p_{1})}}{\sqrt{p_{1}(1-p_{1})}+\sqrt{p_{2}(1-p_{2})}}$ (22)
+Next we want to examine whether for a given $(1,1)$ decomposition with $p_{1}$, $p_{2}$, and $p$ it is convenient to split the pure state into two, having a $(1,2)$ decomposition.
+It can be seen that the ratios $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ are equal with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}_{2}}^{2}}$ (see left panel in Fig.
+4).
+For the new $p^{\prime}_{2}$ of this new state at phase $\phi_{2}$ the calculation gives
+$p^{\prime}_{2}(p_{1},p;\phi_{2})=p_{1}+(p-p_{1})\frac{2p_{1}(1-p_{1})-(p-p_{1})(2p_{1}-1)\cos^{2}\phi_{2}+\cos\phi_{2}\sqrt{4p(1-p)p_{1}(1-p_{1})+(p-p_{1})^{2}\cos^{2}\phi_{2}}}{2(p_{1}(1-p_{1})+(p-p_{1})^{2}\cos^{2}\phi_{2})}$ (23)
+Figure 5: Convexification procedure of $|Z3\rangle$ mixed with the corresponding orthogonal state.
+The curves are made convex with a line touching at $p_{c}=0.0964142$, as is highlighted in the inset, were we plot the difference in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line.
+The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are shown in blue above both curves.
+B.2 $(2,2)$-decompositions
+In the following we will assume that two complex conjugated pairs, corresponding to a superposition of two states at $p_{i}$ and angles $\pm\phi_{i}$, $i\in\{1,2\}$, such that both states come to lie inside the Bloch sphere projection.
+We find
+$m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $\frac{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (24)
+$=$ $\frac{l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)}{l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)+l_{2}(p_{1},\phi_{1};p_{2},\phi_{2};p)}$ (25)
+$m_{1}(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $1-m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ (26)
+$p(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $p_{1}+(p_{2}-p_{1})\frac{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (27)
+$=$ $\frac{p_{2}\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+p_{1}\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (28)
+$=$ $p_{1}m_{1}(p_{1},\phi_{1};p_{2},\phi_{2})+p_{2}m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ (29)
+$l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)$ $=$ $2\sqrt{(p_{1}-p)^{2}+p_{1}(1-p_{1})\cos^{2}\phi_{1}}$ (30)
+$l_{2}(p_{1},\phi_{1};p_{2},\phi_{2};p)$ $=$ $2\sqrt{(p_{2}-p)^{2}+p_{2}(1-p_{2})\cos^{2}\phi_{2}}$ (31)
+From this it is seen that the weights of the states are symmetric under exchanging $p_{i}\to 1-p_{i}$ separately for $i\in\{1,2\}$.
+Setting $\phi_{2}=0$, we are in $(2,1)$ decompositions and the result is
+$p_{2}(p_{1},\phi_{1};p)=p_{1}+(p-p_{1})\frac{(p_{1}-p)(2p_{1}-1)+2p_{1}(1-p_{1})\cos^{2}\phi_{1}+\sqrt{(p_{1}-p)^{2}+4p(1-p)p_{1}(1-p_{1})\cos^{2}\phi_{1}}}{2((p_{1}-p)^{2}+p_{1}(1-p_{1})\cos^{2}\phi_{1})}$ (32)
+Next, for the evaluation of $(2,2)$ from $(2,1)$ decompositions for given $p_{1}$, $\phi_{1}$, and $p$ the ratios of the following quantities are kept fixed (see left panel of Fig.
+4): $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}}_{2}^{2}}$.
+These have to be substituted in $m_{2}(p_{1},\phi_{1};p^{\prime}_{2},\phi^{\prime}_{2})$ where $x^{\prime}_{2}=2\sqrt{p^{\prime}_{2}(1-p^{\prime}_{2})}\cos\phi^{\prime}_{2}$.
+Figure 6: Convexification procedure of the two corresponding pure states lying on the line given by the vector $|Z_{1}\rangle+|Z_{2}\rangle$ in a mixed state.
+The curves are made convex with a line touching at $p_{c}=0.962243$, as is highlighted in the inset, were we plot the difference in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line.
+The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are above both curves; they are not shown here.
+Appendix C The pure entangled states $|N_{i}\rangle$
+We analyze the Bloch sphere of the rank-two mixed state
 $\rho[p]=p|GHZ\rangle\langle GHZ|+(1-p)|W\rangle\langle W|$ (33)
@@
 The one orthogonal decomposition chosen here is the real one of the three pure states of the zero-polytope together with its orthogonal partner.
+Figure 7: Left: Difference of $\sqrt{\tau_{3}}$ of $(1,1)$ and $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
+Right: Positions of the entangled pure state in the $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
+The inset shows the calculated distance of the plane going through three calculated states to the center of the Bloch sphere.
 These are given by
@@
 $|Z_{3;\perp}\rangle$ $=$ $\sqrt{1-p_{0}}|GHZ\rangle+\sqrt{p_{0}}|W\rangle$ (35)
-[…]
-Figure 2: Convexification procedure of $|Z3\rangle$ mixed with the corresponding orthogonal state.
-The curves are made convex with a line touching at $p_{c}=0.0964142$, as is highlighted in the inset, were we plot the differenz in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line.
-The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are shown in blue above both curves.
+with $p_{0}=\frac{4\sqrt[3]{2}}{3+4\sqrt[3]{2}}$ Lohmayer et al.
+(2006).
+The relevant charcteristic curves for $(1,1)$ decompositions with the W state and $(2,1)$ decompositions with $(|Z_{1}\rangle+|Z_{2}\rangle)/\sqrt{2}$ are shown in Fig.
+5 together with their convexification.
 It shows that at $p_{c}=0.0964142$ the line is a proper convexification, starting from zero and touching the curve for $(2,1)$ decompositions; the $(1,1)$ decompositions are shown in blue.
@@
 The angle of the new state is hence calculated to be $\theta=\theta_{0}+\theta_{1}=1.99158=114.109^{\circ}$.
-The four tetrahedra are shown in grey in Fig.
+The four tetrahedra are shown in grey in the right panel of Fig.
 4 with green edges.
@@
 This shows as well that in this case $(2,1)$ decompositions are optimal in the direction of $|GHZ\rangle$ and $(1,1)$ decompositions are optimal in the direction of $|W\rangle$.
-Fig 7 shows these $(2,1)$ decompositions as blue lines on the Bloch sphere surface.
-Figure 3: Convexification procedure of the two corresponding pure states lying on the line given by the vector $|Z_{1}\rangle+|Z_{2}\rangle$ in a mixed state.
-The curves are made convex with a line touching at $p_{c}=0.962243$, as is highlighted in the inset, were we plot the differenz in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line.
-The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are above both curves; they are not shown here.
+The right panel in Fig 4 shows these $(2,1)$ decompositions as blue lines on the Bloch sphere surface.
 In Neveling and Osterloh (2024), much effort has been given concerning the distribution of optimal decompositions.
@@
 The corresponding characteristic curves can be viewed in Fig.
-3.
+6.
 It is seen that a convexification leads here to a $p_{c}=0.962243$ which corresponds to $p_{2}(p_{0},\pi,p_{c})=0.989858$ for the corresponding pure state on the Bloch sphere surface.
@@
 So, the angle of the state $|M_{3}\rangle$ is $\theta=\theta_{0}+\theta_{1}=2.25566=129.24^{\circ}$.
-The three points corresponding to the states $|M_{i}\rangle$ for $i=1,2,3$ are marked in Fig.
+The three points corresponding to the states $|M_{i}\rangle$ for $i=1,2,3$ are marked in the right panel in Fig.
 1 with white dots.
@@
 It is important noticing that, at first sight, the effective tangle of $(2,1)$ and $(1,1)$ decompositions yield the same result.
-That this is not so is seen by looking at the difference between $(1,1)$ and $(2,1)$ decompositions, shown in Fig.
-4.
+That this is not so is seen by looking at the difference between $(1,1)$ and $(2,1)$ decompositions, shown in the left panel of Fig.
+7.
 $\lambda$ refers to the parameter convexly connecting $|N_{3}\rangle$ and $|M_{3}\rangle$, whereas $\lambda_{2}$ is the corresponding parameter for $|W\rangle$ and $|Z_{3}\rangle$.
-Figure 4: Difference of $\sqrt{\tau_{3}}$ of $(1,1)$ and $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
 It is seen that the states under consideration get optimised by choosing the $(2,1)$ decomposition.
 We optimize the tangle with respect to $\lambda_{2}$.
-The pure states are located in the positions indicated in Fig.
-5.
-Figure 5: Positions of the entangled pure state in the $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
-This leads to the lower blue curves shown in Fig.
+The pure states are located in the positions indicated in the right panel of Fig.
+7.
+This leads to the lower blue curves shown in the right panel of Fig.
 4 which lie on a circle that is not a grand circle as the ones before; instead it has a distance to the center of the Bloch sphere of $0.0711148$.
-There are numerical errors to this, but they are small and only occur in the border regions, where the difference to the $(1,1)$ decompositions tends to vanish (see Fig.
-6)
-Figure 6: Difference of $\sqrt{\tau_{3}}$ of $(1,1)$ and $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$.
+This is calculated by planes going through three calculated states to the center of the Bloch sphere.
+In addition to the states $N_{i}$ and $M_{j}$ in between we minimize the tangle on a finite but arbitrary grating.
+We modified the distance the points should have on the grating.
+The result did not depend on this choice.
+The result equals this for the two states $N_{i}$ and the state $M_{j}$ in between.
+There are numerical errors (due to a finite grating) to this, but they are small (smaller than 44‰) and only occur in the border regions, where the difference to the $(1,1)$ decompositions tends to vanish (see inset of the right panel of Fig.
+7).
 That the states form a circle is confirmed by calculating the derivative with respect to the angle and showing its orthogonality to the original vector (not shown here).
 The numerical error scales with $\Delta\phi^{2}$ as expected.
-The normal direction of the circle is $(0.57589,0,-0.81753)$.
+The normal direction of the circle is (0.57589,0,0.81753)(0.57589,0,-0.81753).
 Using the symmetry, we get the three points of normal vectors together with a dashed line in blue.
@@
 For the remaining points in the Bloch sphere there is only a single visible state from the zero polytope left which leads to $(1,1)$ decomposition being optimal there.
-They are depicted in Fig.
+They are depicted in the right panel of Fig.
 4 by green lines and red points.
-Figure 7: Optimal decompositions are obtained for the rank-two mixtures of GHZ and W state and are depicted here graphically.
-It consists of the zero polytope (blue shaded), 4 polytopes defined by three of the zero states together with a state $|N\rangle$ one of which is the GHZ state itself (grey shaded); all these polytopes are three dimensional.
-On top of this kernel structure of five polytopes are three lines along grand circles along $\phi$ direction and three lines which are on circles with a distance $0.0711148$ to the center of the Bloch sphere.
-Along these lines move pure states which form a two-dimensional optimal decomposition (red triangles).
-Blue points mark the position of the respective normal vectors on the Bloch sphere surface; blue lines the corresponding pure states in $(2,1)$ decompositions.
-All remaining parts of the Bloch sphere are covered by $(1,1)$ decompositions.
-Conclusions – The mixture of GHZ and W state has been elaborated for all the Bloch sphere, obtaining a pattern on the surface that is characteristic for these two states in the mixture.
-The finding is based on the zero-state locking and that $(n_{z},n_{e})$ decompositions for $n_{e}>1$ are not optimal for the present case.
-We have checked this numerically but there is no general proof for that finding.
-For the zero-state locking it is essential to look at a certain roots of the tangle, which is the square root in the case of the threetangle.
-The Bloch sphere is divided into the zero polytope and four polytopes onto each faces of the zero polytope with corresponding four pure states.
-The faces of two neighboring such polytopes are strictly planar.
-There are pure states that belong to $(2,1)$ decompositions moving on three azimuthal grand circles and three circles which possess a small distance $0.0711148$ to the Bloch sphere center.
-Their normal vector is $(0.57589\cos\phi_{i},0,-0.81753\sin\phi_{i})$, with $\phi_{i}=i2\pi/3$.
-The remaining decompositions to fill the Bloch sphere are of the type $(1,1)$.
-This distribution of the optimal decomposition is natural and inter-connects the polytopes of various dimensions.
-For SL invariant tangles there is a transformation rule that applies in particular to the optimal decomposition statesViehmann et al.
-(2012).
-Applying these rules yields the solution for the convex-roof for these cases.
-For the general case the characteristic features of the convex roof will modify, but the structure will stay the same, as reported also in ref.
-Neveling and Osterloh (2024).
-It would be interesting to what extent the findings are universal.
-The paths of the pure entangled state of $(2,1)$ decompositions here moves on circles, but this does not need to be this way.
-Future investigation, clarifying this question, will be needed.
-The analysis is based on two essential pillars: the zero-state locking and that every optimal decomposition would be of the type $(n_{z},1)$.
-It would be crucial to either prove or disprove these conjectures in future work.
-This presumably clarifies how optimal decompositions of rank two mixtures behave. It will be very intriguing to understand their behavior for higher ranks.
-Acknowledgements – I acknowledge the TII and Luigi Amico for supporting this research.
-Appendix A General properties of optimal decompositions for real states of rank two
-For the number $n_{\rm opt}$ of pure states contained in any optimal decomposition holds ${\rm rank}[\rho]\leq n_{\rm opt}\leq({\rm rank}[\rho])^{2}$ Carathéodory (1911); Uhlmann (1998).
-Thus, it is sufficient to look for up to $4$ such pure states that generically form three-dimensional simplices.
-We want to emphasize that more than the maximal $4$ states may be in an optimal decomposition; in this case, since every sub-partition of an optimal decomposition is itself optimal, each $4$ states out of that decomposition are optimal as well and the optimal decomposition is given by the convex polytope made out of these points.
-The tangle will behave linearly in these optimal polytopes.
-Since the states are composed of real elements every tangle $\tau$ has the property that $\tau(z)\equiv\tau(z^{*})$ where $|\psi\rangle\propto|\psi_{1}\rangle+z|\psi_{2}\rangle$ is a representation that corresponds to a vector onto the Bloch-sphere with the parametrizationOsterloh (2016)
-$z\mathrel{\widehat{=}}\vec{n}=\left(\begin{array}[]{c}x\\ y\\ z\end{array}\right)=\left(\begin{array}[]{c}2\sqrt{p(1-p)}\cos\phi\\ 2\sqrt{p(1-p)}\sin\phi\\ 2p-1\end{array}\right)$ (4)
-where $|GHZ\rangle$ and $|W\rangle$ are the two anti-podes on the z-axis. Real superpositions lie in the $x$-$z$-plane.
-We will be mainly interest in a) where the line $\overline{\vec{n}_{0}\vec{n}_{1}}$ cuts the $z$-axis, b) which are the weights of the corresponding entangled state(s), and c) where it crosses the Bloch sphere, that is the corresponding pure states.
-Whereas the answer to a) is given by the probability
-$P=\frac{p_{0}\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}+p_{1}\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{p_{0}\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}+p_{1}\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (5)
-such that the point is located in $\vec{n}_{P}=(2P-1)\vec{e}_{z}$, the answer to b) is given by
-$\lambda=\frac{\mu_{1}}{\mu_{0}}=\frac{\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (6)
-and for the mainly interesting case of $\phi_{1}=0$ and $p$ being the probability in the density matrix in consideration, we obtain as a solution to c)
-$P_{\pm}=p_{0}+\frac{p-p_{0}}{2}\frac{(p-p_{0})(1-2p_{0})+2p_{0}(1-p_{0})\cos^{2}\phi_{0}\pm\sqrt{(p_{0}-p)^{2}+4pp_{0}(1-p_{0})(1-p)\cos^{2}\phi_{0}})}{(p-p_{0})^{2}+p_{0}(1-p_{0})\cos^{2}\phi_{0}}$ (7)
-The resulting tangle is a convex combination of the tangle values of the respective pure states.
-Figure 8: States within the Bloch sphere.
-For a given line $\overline{x_{1}x_{2}}$ all triangles $z_{0}x_{1}z_{1}$, $z_{0}x_{2}z_{2}$, and $z_{0}x_{2}^{\prime}z_{2}^{\prime}$, are similar.
-Ratios of their length are determined by the theorem of intersecting lines.
-This strictly linear behavior inside all the simplices of an optimal decomposition indirectly tells about whether there needs to be one or more states in a decomposition for being optimal: the continuation of the respective decomposition type has to lead to a convex behavior in the tangle value; where this condition is not satisfied a (linear) convexification is needed.
-This holds in particular if the corresponding optimal polytope changes dimension.
-Here, this means to a dimensionality of at most three.
-A.1 $(1,1)$-decompositions
-Here, we collect the formulae obtained for $(1,1)$-decompositions of a pure state at z-coordinate $12p_{1}-$ 1, where $p_{1}$ corresponds to its probability of the two states $|\psi_{i}\rangle$; $i\in\{N,S\}$.
-This corresponds to two pure states in the Fig.
-8.
-We find
-$l_{1}(p_{1},p)$ $=$ $2\sqrt{(p-p_{1})^{2}+p_{1}(p_{1}-1)}$ (8)
-$l_{2}(p_{1},p)$ $=$ $\frac{4p(1-p)}{l_{1}(p_{1},p)}$ (9)
-$m_{2}(p_{1},p)$ $=$ $\frac{l_{1}(p_{1},p)}{l_{1}(p_{1},p)+l_{2}(p_{1},p)}$ (10)
-$p_{2}(p_{1},p)$ $=$ $\frac{4p^{2}(1-p_{1})}{l_{1}(p_{1},p)^{2}}$ (11)
-$=$ $(1-p_{1})\frac{p}{1-p}\cdot\frac{l_{2}(p_{1},p)}{l_{1}(p_{1},p)}$ (12)
-$p(p1,p2)$ $=$ $p_{1}+\frac{(p_{2}-p_{1})\sqrt{p_{1}(1-p_{1})}}{\sqrt{p_{1}(1-p_{1})}+\sqrt{p_{2}(1-p_{2})}}$ (13)
-Next we want to examine whether for a given $(1,1)$ decomposition with $p_{1}$, $p_{2}$, and $p$ it is convenient to split the pure state into two, having a $(1,2)$ decomposition.
-It can be seen that the ratios $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ are equal with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}_{2}}^{2}}$ (see Fig.
-8).
-For the new $p^{\prime}_{2}$ of this new state at phase $\phi_{2}$ the calculation gives
-$p^{\prime}_{2}(p_{1},p;\phi_{2})=p_{1}+(p-p_{1})\frac{2p_{1}(1-p_{1})-(p-p_{1})(2p_{1}-1)\cos^{2}\phi_{2}+\cos\phi_{2}\sqrt{4p(1-p)p_{1}(1-p_{1})+(p-p_{1})^{2}\cos^{2}\phi_{2}}}{2(p_{1}(1-p_{1})+(p-p_{1})^{2}\cos^{2}\phi_{2})}$ (14)
-A.2 $(2,2)$-decompositions
-In the following we will assume that two complex conjugated pairs, corresponding to a superposition of two states at $p_{i}$ and angles ±$\pm\phi_{i}$, $i\in\{1,2\}$, such that both states come to lie inside the Bloch sphere projection.
-We find
-$m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $\frac{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (15)
-$=$ $\frac{l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)}{l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)+l_{2}(p_{1},\phi_{1};p_{2},\phi_{2};p)}$ (16)
-$m_{1}(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $1-m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ (17)
-$p(p_{1},\phi_{1};p_{2},\phi_{2})$ $=$ $p_{1}+(p_{2}-p_{1})\frac{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (18)
-$=$ $\frac{p_{2}\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+p_{1}\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}{\sqrt{p_{1}(1-p_{1})\cos^{2}(\phi_{1})}+\sqrt{p_{2}(1-p_{2})\cos^{2}(\phi_{2})}}$ (19)
-$=$ $p_{1}m_{1}(p_{1},\phi_{1};p_{2},\phi_{2})+p_{2}m_{2}(p_{1},\phi_{1};p_{2},\phi_{2})$ (20)
-$l_{1}(p_{1},\phi_{1};p_{2},\phi_{2};p)$ $=$ $2\sqrt{(p_{1}-p)^{2}+p_{1}(1-p_{1})\cos^{2}\phi_{1}}$ (21)
-$l_{2}(p_{1},\phi_{1};p_{2},\phi_{2};p)$ $=$ $2\sqrt{(p_{2}-p)^{2}+p_{2}(1-p_{2})\cos^{2}\phi_{2}}$ (22)
-From this it is seen that the weights of the states are symmetric under exchanging $p_{i}\to 1-p_{i}$ separately for $i\in\{1,2\}$.
-Setting $\phi_{2}=0$, we are in $(2,1)$ decompositions and the result is
-$p_{2}(p_{1},\phi_{1};p)=p_{1}+(p-p_{1})\frac{(p_{1}-p)(2p_{1}-1)+2p_{1}(1-p_{1})\cos^{2}\phi_{1}+\sqrt{(p_{1}-p)^{2}+4p(1-p)p_{1}(1-p_{1})\cos^{2}\phi_{1}}}{2((p_{1}-p)^{2}+p_{1}(1-p_{1})\cos^{2}\phi_{1})}$ (23)
-Next, for the evaluation of $(2,2)$ from $(2,1)$ decompositions for given $p_{1}$, $\phi_{1}$, and $p$ the ratios of the following quantities are kept fixed (see Fig.
-8): $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}}_{2}^{2}}$.
-These have to be substituted in $m_{2}(p_{1},\phi_{1};p^{\prime}_{2},\phi^{\prime}_{2})$ where $x^{\prime}_{2}=2\sqrt{p^{\prime}_{2}(1-p^{\prime}_{2})}\cos\phi^{\prime}_{2}$.